# Multi-scale Latent Point Consistency Models for 3D Shape Generation

### v1

@@ Abstract
Consistency Models (CMs) have significantly accelerated the sampling process in diffusion models, yielding impressive results in synthesizing high-resolution images. To explore and extend these advancements to point-cloud-based 3D shape generation, we propose a novel Multi-scale Latent Point Consistency Model (MLPCM). Our MLPCM follows a latent diffusion framework and introduces hierarchical levels of latent representations, ranging from point-level to super-point levels, each corresponding to a different spatial resolution. We design a multi-scale latent integration module along with 3D spatial attention to effectively denoise the point-level latent representations conditioned on those from multiple super-point levels. Additionally, we propose a latent consistency model, learned through consistency distillation, that compresses the prior into a one-step generator. This significantly improves sampling efficiency while preserving the performance of the original teacher model. Extensive experiments on standard benchmarks ShapeNet and ShapeNet-Vol demonstrate that MLPCM achieves a 100x speedup in the generation process, while surpassing state-of-the-art diffusion models in terms of both shape quality and diversity.
1 Introduction
@@
Generative modeling of 3D shapes plays a crucial role in various applications in 3D computer vision and graphics, empowering digital artists to create realistic, high-quality shapes. For these models to be practically effective, they must provide flexibility for interactive refinement, support the synthesis of diverse shape variations, and generate smooth meshes that seamlessly integrate into standard graphics pipelines. With the rapid advances in generative models for text, images, and videos, significant progress has also been made in 3D shape generation. Techniques based on variational autoencoders (VAEs) [22, 51, 33], generative adversarial networks (GANs)  [54, 44, 8, 12], and normalizing flow models [55, 13, 18] have been proposed to generate 3D shapes. Recently, diffusion-based models [34] and their latent diffusion variants [56, 41] have achieved state-of-the-art results in this domain.
Despite these advancements, challenges remain with applying diffusion models to 3D shapes represented as point clouds. First, the irregular spatial distribution and large number of points require the denoising model to capture both local and global geometric patterns accurately to produce high-quality 3D shapes. Diffusing in the latent space, where 3D shapes are summarized into more compact representations, is arguably more appealing than operating directly in data space, as it simplifies the learning process. However, prior work has demonstrated that relying on a single level (scale or spatial resolution) of latent representations is insufficient for achieving satisfactory performance. How to effectively leverage information from multiple levels remains a critical question. Second, the sampling process in diffusion models is notoriously slow, making them impractical for real-world applications.
In this paper, we introduce a novel model, Multi-scale Latent Points Consistency Models (MLPCM), which follows the latent diffusion framework. Our approach is a hierarchical VAE that incorporates multiple levels of latent representations, ranging from point-level to super-point levels, each corresponding to a different scale (spatial resolution). We establish a hierarchical dependency among these multi-scale latent representations in the VAE encoder. Diffusion is then performed on the point-level latent representations, where noisy point-level latent inputs are modulated by latent representations from multiple super-point levels through a specially designed multi-scale latent integration module. This design allows our model to better capture both local and global geometric patterns of 3D shapes. To further improve performance, we incorporate a 3D spatial attention mechanism into the denoising function, where the bias term is determined by pairwise 3D distances, encouraging the model to focus on nearby regions within the 3D space. After training the model, we leverage consistency distillation to learn a latent consistency model, which significantly accelerates the sampling process.
@@
  • We propose a multi-scale latent integration module along with the 3D spatial attention mechanism for effectively improving the denoising process in the latent point space.
  • We explore the latent consistency distillation to achieve fast (one-step) and high-quality 3D shape generation.
@@
2 RELATED WORKS
Diffusion Models.
Diffusion models have seen significant success in image generation [9, 48, 49, 47, 40, 42, 35]. These models are trained to denoise data that has been corrupted by noise, thereby estimating the score of the data distribution. During inference, they generate samples by running the reverse diffusion process, which gradually removes noise from the data points. Compared to VAEs [16, 45] and GANs [7], diffusion models offer advantages in terms of training stability and more accurate likelihood estimation.
Accelerating DMs.
Despite their success, diffusion models are limited by slow generation speeds. To address this, various approaches have been proposed. Training-free methods include ODE solvers [48, 28, 29], adaptive step-size solvers [11], and predictor-corrector methods [48]. Training-based strategies involve optimized discretization [53], truncated diffusion [31, 59] , neural operators [60], and distillation techniques [43, 32]. Additionally, more recent generative models have been introduced to enable faster sampling [24, 25]. [50] have demonstrated that Consistency Models (CMs) hold great promise as a new type of generative model, offering faster sampling while maintaining high-quality generation. CMs utilize consistency mapping to directly map any point along an ODE trajectory back to its origin, allowing for rapid one-step generation. These models can be trained either by distilling pre-trained diffusion models or as independent generative models. Further details about CMs are provided in the following section.
3D Point Cloud Generation.
@@ 2 RELATED WORKS
This task is also referred to as shape generation. Achlioptas et al. [1] were were the first to study this problem, developing a straightforward GAN with multiple MLPs in both the generator and discriminator. They also introduced several metrics to evaluate the quality of 3D GANs. Valsesia et al. [52] improved the generator by incorporating graph convolution operations. Liu et al. [23] employed a tree-structured graph convolution network (GCN) to capture the hierarchical information of parent nodes. Gal et al. [6] uses a GAN with multiple roots to generate point sets that achieve unsupervised part disentanglement.
@@
Beyond GAN-based methods, there are also approaches rooted in probability theory [57, 21, 5]. ShapeGF Cai et al. [3] uses stochastic gradient ascent on an unnormalized probability density to move randomly sampled points toward high-density regions near the surface of a specific shape. DPM [30] conceptualized the transformation of points from a noise distribution to a point cloud as the inverse process of particle diffusion in a thermal system in contact with a heat bath. They modeled this inverse diffusion as a Markov chain conditioned on a particular shape. [55] introduced PointFlow, which generates point sets by modeling them as a distribution of distributions within a probabilistic framework. [15] presented a flow-based model called ChartPointFlow, which constructs a map conditioned on a label, preserving the shape’s topological structure. [20] developed a modified VAE for parts-aware editing and unsupervised point cloud generation.
@@
3 Multi-scale Latent Point Diffusion Models and Latent Consistency Models
Figure 1: An overview of the proposed multi-scale latent diffusion model. Both the VAE and the latent point diffusion model use Set Abstraction modules and Feature Propagation modules. The details of the diffusion prior are shown in Figure 2.
@@ 3 Multi-scale Latent Point Diffusion Models and Latent Consistency Models
3.1 Hierarchical VAE Framework
We begin by formally introducing the latent diffusion framework, which is essentially a hierarchical VAE. We denote a point cloud as $X\in\mathbb{R}^{N\times 3}$, consisting of $N$ points with 3D coordinates. We then introduce a hierarchy of latent variables with different spatial scales/resolutions, denoting as $\mathcal{Z}=\{Z^{0},Z^{1},\dots,Z^{L}\}$, where $L$ is the number of hierarchy levels. We use superscript to denote the index of scales. Here $Z^{l}\in\mathbb{R}^{N_{l}\times C_{l}}$ where $N_{l}$ and $C_{l}$ are the number of points and the number of channels at the $l$-th level latent space respectively. Note that $C_{l}=3+D_{l}$, where $D_{l}$ represents the extra dimension. At the 00-th, we have $N_{0}=N$ and $N_{l}>N_{l+1}$ for any $l=0,\dots,L-1$. In other words, the $Z^{0}$ is latent point representation whereas $Z^{l}$ are latent super-point (i.e., a subset of points) representations for any $l>0$. The details of latent layers and points in each level are demonstrated in Section 4 .
@@ 3.1 Hierarchical VAE Framework
$-\lambda D_{KL}(q_{\phi}(\mathcal{Z}|X)\|p(\mathcal{Z}))]$ (2)
@@
Figure 2: An illustration of the network architecture of the latent diffusion prior. We denote the voxel grid size as $r$, and the hidden dimension as $D$ and $D^{{}^{\prime}}$.
3.2 Multi-Scale Latent Point Diffusion As Prior
@@
$p_{\theta}(Z^{0}_{0:T}|\mathcal{Z}^{\setminus{0}})$ $:=p(Z^{0}_{T})\prod_{t=1}^{T}p_{\theta}(Z^{0}_{t-1}|Z^{0}_{t},\mathcal{Z}^{\setminus{0}})$ (6)
@@
3.3 Architecture Design
@@
[…]
@@
$F_{v}=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d}}+B\right)\cdot V$ (10)
@@
3.4 Multi-Scale Latent Point Consistency Models
@@
$\frac{\mathrm{d}Z^{0}_{t}}{\mathrm{d}t}=\frac{\mathrm{d}\log\alpha_{t}}{\mathrm{d}t}Z^{0}_{t}+\left(\frac{\mathrm{d}\log\alpha_{t}}{2\sigma_{t}\mathrm{d}t}-\frac{\mathrm{d}\log\alpha_{t}}{\mathrm{d}t}\sigma_{t}\right)\epsilon_{\theta}(Z^{0}_{t},t,\mathcal{Z}^{\setminus{0}})$ (11)
where $\epsilon_{\theta}(Z^{0}_{t},t,\mathcal{Z}^{\setminus{0}})$ is the noise prediction model and $\alpha_{t},\sigma_{t}$ are determined by the noise schedule as aforementioned. Samples can be drawn by solving the PF-ODE from $T$ to 00. To achieve consistency distillation, we introduce a consistency map $f_{\hat{\theta}}:(Z^{0}_{t},t,\mathcal{Z}^{\setminus{0}})\mapsto Z_{0}^{0}$ to directly predict the solution of PF-ODE in Equation 11. For $t=0$, we parameterize $f_{\hat{\theta}}$ using the noise prediction model $\hat{\epsilon}_{\theta}$:
$f_{\hat{\theta}}(Z^{0}_{t},t,\mathcal{Z}^{\setminus{0}})=c_{\text{skip}}(t)Z^{0}_{t}+c_{\text{out}}(t)\left(\frac{Z^{0}_{t}-\sigma_{t}\hat{\epsilon}_{\hat{\theta}}(Z^{0}_{t},t,\mathcal{Z}^{\setminus{0}})}{\alpha_{t}}\right),$ (12)
@@
$\mathcal{L}_{\text{CD}}(\hat{\theta},\hat{\theta}^{-};\Psi)=$
$\mathbb{E}_{n,\mathcal{Z}^{\setminus{0}},Z^{0}_{t_{n+1}}}\left[d\left(f_{\hat{\theta}}\left(Z^{0}_{t_{n+1}},\mathcal{Z}^{\setminus{0}},t_{n+1}\right),f_{\hat{\theta}^{-}}\left(\hat{Z}^{0}_{t_{n}},\mathcal{Z}^{\setminus{0}},t_{n}\right)\right)\right],$ (13)
where $\hat{Z}^{0}_{t_{n}}=Z^{0}_{t_{n+1}}+(t_{n}-t_{n+1})\Psi(Z^{0}_{t_{n+1}},t_{n+1},t_{n},\mathcal{Z}^{\setminus{0}})$ is the solution of the PF-ODE obtained via calling the ODE solver $\Psi$ from time $t_{n+1}$ to $t_{n}$. $d(\bm{x},\bm{y})=\|\bm{x}-\bm{y}\|^{2}_{2}$.
4 Experiments

### v2

@@ Abstract
Consistency Models (CMs) have significantly accelerated the sampling process in diffusion models, yielding impressive results in synthesizing high-resolution images. To explore and extend these advancements to point-cloud-based 3D shape generation, we propose a novel Multi-scale Latent Point Consistency Model (MLPCM). Our MLPCM follows a latent diffusion framework and introduces a hierarchy of latent representations, ranging from point-level to super-point levels, each corresponding to a different spatial resolution. We design a multi-scale latent integration module along with 3D spatial attention to effectively denoise the point-level latent representations conditioned on those from multiple super-point levels. Additionally, we propose a latent consistency model, learned through consistency distillation, that compresses the prior into a one-step generator. This significantly improves sampling efficiency while preserving the performance of the original teacher model. Extensive experiments on standard benchmarks ShapeNet and ShapeNet-Vol demonstrate that MLPCM achieves a 100x speedup in the generation process, while surpassing state-of-the-art diffusion models in terms of both shape quality and diversity.
I Introduction
@@
Generative modeling of 3D shapes plays a crucial role in various applications in 3D computer vision and graphics, empowering digital artists to create realistic and high-quality shapes [1, 2, 3, 4, 5, 6, 7]. For these models to be practically effective, they must provide flexibility for interactive refinement, support the synthesis of diverse shape variations, and generate smooth meshes that seamlessly integrate into standard graphics pipelines.
With the rapid advances in generative models for text, images, and videos, significant progress has also been made in 3D shape generation. Techniques based on variational autoencoders (VAEs) [8, 9, 10], generative adversarial networks (GANs)  [11, 12, 13, 14], and normalizing flow models [2, 15, 16] have been proposed to generate 3D shapes. Recently, diffusion-based models [17] and their latent diffusion variants [3, 18] have achieved state-of-the-art results in this domain.
Despite notable progress, substantial obstacles remain when applying diffusion models to 3D shapes that are represented as point clouds. Point sets are unordered, irregular, and often highly non-uniform in density, which complicates both neighborhood construction and feature aggregation [19, 20]. A single radius may fail to capture thin structures while a large radius may blur sharp edges, and k-nearest-neighbor graphs can become unstable under varying sampling density [20, 21]. Real data frequently contain noise, outliers, holes, and partial coverage from limited viewpoints [22, 23], so the denoiser must be robust to missing regions while still reconstructing fine details. In addition, 3D geometry introduces invariances and symmetries that are easy to violate. Models that are not carefully designed with respect to translation, rotation, and scale may overfit to coordinate frames or camera poses [24, 25]. These factors push the denoising network to learn representations that respect both local geometric primitives such as edges, corners, and small parts, and global object level regularities such as symmetry, topology, and part arrangement. Capturing these patterns jointly is essential for producing high quality point sets that are faithful at a small scale while remaining structurally coherent at a large scale.
A promising strategy is to diffuse in a learned latent space where shapes are summarized into compact codes that suppress nuisance variation and concentrate information about structure [26, 3]. Learning in latent space reduces computational cost and can improve stability, yet relying on a single level of latent representation has proved insufficient. A single scale cannot simultaneously preserve tiny geometric details and encode global layout, which leads to oversmoothed reconstructions or broken long range dependencies [27, 28]. Effective 3D generation therefore requires multiple levels or resolutions together with mechanisms that let information flow across them so that coarse latents guide overall structure while finer latents supply detail where needed [3]. A second challenge concerns practicality. The sampling process in diffusion models typically requires many iterative evaluations of the denoiser, which results in slow generation and high memory consumption [29]. Classifier free guidance and other conditioning signals further increase the computational burden [30]. Without acceleration, these costs hinder real world use in interactive modeling, editing, completion from sparse inputs, or deployment within larger 3D pipelines [31, 32]. Reducing the number of steps while preserving quality, for example through consistency distillation or higher order solvers in latent space [33, 34], remains critical for making diffusion-based point cloud generators both accurate and efficient.
Here we introduce Multi-scale Latent Point Consistency Models (MLPCM), a generative model for 3D shape synthesis that follows the latent diffusion paradigm while addressing the joint learning of fine local geometry and long range structural coherence within a single framework. The architecture is a hierarchical variational autoencoder that encodes each shape into multiple levels of latent variables, from point level to super-point levels, each aligned with a specific spatial resolution. These latents form an explicit hierarchy in the encoder so that information flows in both top down and bottom up directions. Coarse semantic latents provide guidance for the organization of global structure, while fine latents inject detail that refines and corrects coarse hypotheses. Diffusion operates on point level latents, which concentrates stochasticity where geometric detail resides and keeps higher level latents clean and semantically stable. A multi-scale latent integration module then conditions the denoiser by modulating noisy point latents with super-point latents from several scales. This cross-scale conditioning encourages the emergence of compositional regularities, supports reuse of part-level patterns across categories, and improves data efficiency under sparse or partial observations.
To further strengthen geometric fidelity and controllability, the denoising network integrates a 3D spatial attention mechanism whose bias term derives from pairwise Euclidean distances. This bias respects rigid transformations and encourages attention to concentrate on geometrically proximate regions while preserving the capacity to aggregate information across the entire shape when long range dependencies are beneficial. Training optimizes a hierarchical evidence lower bound for the VAE together with the standard latent diffusion objective. After training, consistency distillation in latent space yields a latent consistency model that preserves the visual quality of the diffusion sampler while reducing the number of refinement steps, which enables rapid sampling for interactive editing and conditional generation. In combination, hierarchical latent organization, multi-scale integration, geometry aware attention, and accelerated inference provide a practical foundation for downstream tasks such as completion from sparse views, structure-preserving editing, controllable synthesis with part level constraints, and conditional reconstruction from partial point clouds, while also improving robustness to noise, distribution shift, and limited supervision.
@@
  • We propose a multi-scale latent integration module along with a 3D spatial attention mechanism to enable the denoiser to focus on salient neighbourhoods while maintaining global consistency, resulting in sharper topology and more uniform point distribution.
  • We extend consistency distillation to unstructured point‐latent spaces and couple it with multi-scale latent integration, yielding a student that inherits multi-scale semantics and generates 3D shapes in one to four steps without loss of visual or structural fidelity.
@@
II RELATED WORKS
Figure 1: An overview of the proposed multi-scale latent diffusion model. Hierarchical VAEs encode the input shape into latent variables at multiple resolutions, including point-level and super-point-level. Compared to the latent point diffusion model that only denoises the point-level latent space, our proposed multi-scale latent integration module injects higher-level latent variables to guide refinement, thereby jointly modeling local geometry and global structure. The details of the diffusion prior are shown in Figure 2.
In this section, we discuss related works on diffusion models, accelerating diffusion models and 3D point cloud generation, respectively.
II-A Diffusion Models
Diffusion models have achieved state-of-the-art results in image synthesis by learning to reverse a corruption process that gradually adds noise to data. In the common discrete-time formulation, the forward process is a fixed Markov chain that diffuses a clean sample into near-Gaussian noise, while the model is trained to predict either the noise or the score of the noised sample so that it can iteratively denoise at test time [29]. A complementary continuous-time view casts diffusion as stochastic differential equations and learns the score function of perturbed data, which unifies score-based generative modeling with diffusion under a single framework and enables alternative samplers [35, 36]. Maximum likelihood training for continuous-time score models further improves statistical efficiency and connects diffusion training to classic likelihood principles [37].
Scaling laws and architectural choices have made diffusion a robust default for high-fidelity image generation. Hierarchical or two-stage pipelines that combine powerful text encoders with diffusion backbones yield strong semantic control and photorealism [38, 39]. Learning the generative process in a compressed latent space rather than pixel space reduces computational burden and memory cost while maintaining visual quality, which has been key to high-resolution synthesis [26]. Compared with variational autoencoders and generative adversarial networks, diffusion models typically provide more stable optimization, reliable likelihood estimation, and improved mode coverage, mitigating common issues such as posterior collapse in VAEs and mode dropping in GANs [40, 41, 42]. During inference, samples are produced by integrating the learned reverse dynamics from noise toward data, which can be interpreted either as reversing a discrete diffusion chain or integrating an ODE or SDE defined by the learned score [29, 36].
II-B Accelerating Diffusion Models
Despite their success, diffusion models are constrained by slow sampling due to many iterative denoising steps. Training-free acceleration exploits the continuous-time formulation to integrate probability-flow ODE trajectories in fewer steps, including high-order solvers such as DPM-Solver and its improved variants, as well as unified predictor–corrector samplers and adaptive step-size schemes [32, 43, 44]. Training-based acceleration modifies objectives or adds light retraining to shorten trajectories: optimized discretization learns sampling times tailored to a fixed compute budget [45], truncated diffusion shortens the denoising horizon [46, 47], and neural-operator style learners amortize parts of the reverse dynamics [48]. Distillation compresses multi-step teachers into fast students by matching trajectories or terminal outputs, yielding few-step or even single-step generators [49, 50].
Recent works push one-step and few-step generation further. Consistency Models enforce a mapping that is consistent along time and can be trained either by distillation or from scratch, enabling rapid one-step sampling [33]. Follow-up advances improve training stability and data-only training, reduce reliance on perceptual losses, and formalize the theory of consistency training [51, 52]. Stochastic consistency distillation introduces noise-aware consistency to better align with the teacher’s trajectories and improve sample quality at low NFE [53]. In parallel, rectified-flow training has been strengthened to compete with distillation in the low-step regime, showing that a single reflow iteration with improved timestep curricula and robust pre-metrics is often sufficient [54]. Flow-matching and flow trajectory distillation style approaches provide another route to few-step or one-step sampling by directly parameterizing ODE paths from noise to data [55, 56]. There is also growing interest in faster training of diffusion and consistency models themselves, which leverages cross-init consistency phenomena and curriculum-style schedules to cut training cost while retaining generation quality [57]. Together, these strategies reduce latency by one to two orders of magnitude, making diffusion practical for interactive applications and deployment under strict compute budgets.
II-C 3D Point Cloud Generation
@@ II-B Accelerating Diffusion Models
This task is also referred to as shape generation. Achlioptas et al. [58] first proposed to study this problem, developing a straightforward GAN with multiple MLPs in both the generator and discriminator. They also introduced several metrics to evaluate the quality of 3D GANs. Valsesia et al. [59] improved the generator by incorporating graph convolution operations. Liu et al. [60] employed a tree-structured graph convolution network (GCN) to capture the hierarchical information of parent nodes. Gal et al. [61] used a GAN with multiple roots to generate point sets that achieve unsupervised part disentanglement.
@@
Beyond GAN-based methods, there are also approaches rooted in probability theory [62, 63, 64]. ShapeGF [65] used stochastic gradient ascent on an unnormalized probability density to move randomly sampled points toward high-density regions near the surface of a specific shape. DPM [1] conceptualized the transformation of points from a noise distribution to a point cloud as the inverse process of particle diffusion in a thermal system in contact with a heat bath. They modeled this inverse diffusion as a Markov chain conditioned on a particular shape. Yang et al. [2] introduced PointFlow, which generates point sets by modeling them as a distribution of distributions within a probabilistic framework. Kimura et al. [66] presented a flow-based model called ChartPointFlow, which constructs a map conditioned on a label, preserving the shape’s topological structure. Li et al. [67] developed a modified VAE for parts-aware editing and unsupervised point cloud generation.
@@
III Multi-scale Latent Point Diffusion Models and Latent Consistency Models
@@
III-A Hierarchical VAE Framework
We begin by formally introducing the latent diffusion framework, which is essentially a hierarchical VAE. We denote a point cloud as $X\in\mathbb{R}^{N\times 3}$, consisting of $N$ points with 3D coordinates. We then introduce a hierarchy of latent variables with different spatial scales/resolutions, denoting as $\mathcal{Z}=\{Z^{0},Z^{1},\dots,Z^{L}\}$, where $L$ is the number of hierarchy levels. We use superscript to denote the index of scales. Here $Z^{l}\in\mathbb{R}^{N_{l}\times C_{l}}$ where $N_{l}$ and $C_{l}$ are the number of points and the number of channels at the $l$-th level latent space respectively. Note that $C_{l}=3+D_{l}$, where $D_{l}$ represents the extra dimension. At the 0-th, we have $N_{0}=N$ and $N_{l}>N_{l+1}$ for any $l=0,\dots,L-1$. In other words, the $Z^{0}$ is latent point representation whereas $Z^{l}$ are latent super-point (i.e., a subset of points) representations for any $l>0$. The details of latent layers and points in each level are demonstrated in Section IV.
@@ III-A Hierarchical VAE Framework
$-\lambda D_{KL}(q_{\phi}(\mathcal{Z}|X)\|p(\mathcal{Z}))].$ (2)
@@
III-B Multi-Scale Latent Point Diffusion As Prior
@@
$p_{\theta}(Z^{0}_{0:T}|\mathcal{Z}^{\setminus{0}})$ $:=p(Z^{0}_{T})\prod_{t=1}^{T}p_{\theta}(Z^{0}_{t-1}|Z^{0}_{t},\mathcal{Z}^{\setminus{0}}),$ (6)
@@
Figure 2: An illustration of the network architecture of the latent diffusion prior. The multi-scale latent integration module injects higher-level latents to couple local geometry with global structure, while the set abstraction module downsamples and aggregates across scales. We denote the voxel grid size as $r$, and the hidden dimension as $D$.
III-C Architecture Design
@@
where Norm is the layer normalization [69] and Linear denotes a linear layer. MLP consists of two linear layers. $\odot$ denotes element-wise product and $\circ$ means function composition. $\cdot\|\cdot$] means concatenation of vectors. Pos()($\cdot$) is the positional embedding function. To enable scale-time awareness in the model, we further modulate the feature with positional embedding of scale and time. The feature is fed to the following SA and FP modules to predict the noise.
@@
$F_{a}=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d}}+B\right)\cdot V,$ (10)
@@
III-D Multi-Scale Latent Point Consistency Models
@@
$\begin{aligned} \frac{\mathrm{d}Z^{0}_{t}}{\mathrm{d}t}=\frac{\mathrm{d}\log\alpha_{t}}{\mathrm{d}t}Z^{0}_{t}+\left(\frac{\mathrm{d}\log\alpha_{t}}{2\sigma_{t}\mathrm{d}t}-\frac{\mathrm{d}\log\alpha_{t}}{\mathrm{d}t}\sigma_{t}\right)\epsilon_{\theta}(Z^{0}_{t},t,\mathcal{Z}^{\setminus{0}}),\end{aligned}$
where $\epsilon_{\theta}(Z^{0}_{t},t,\mathcal{Z}^{\setminus{0}})$ is the noise prediction model and $\alpha_{t},\sigma_{t}$ are determined by the noise schedule as aforementioned. Samples can be drawn by solving the PF-ODE from $T$ to 0. To achieve consistency distillation, we intrsoduce a consistency map $f_{\hat{\theta}}:(Z^{0}_{t},t,\mathcal{Z}^{\setminus{0}})\mapsto Z_{0}^{0}$ to directly predict the solution of PF-ODE in Equation 11. For $t=0$, we parameterize $f_{\hat{\theta}}$ using the noise prediction model $\hat{\epsilon}_{\theta}$:
$\begin{aligned} f_{\hat{\theta}}(Z^{0}_{t},t,\mathcal{Z}^{\setminus{0}})=c_{\text{skip}}(t)Z^{0}_{t}+c_{\text{out}}(t)\left(\frac{Z^{0}_{t}-\sigma_{t}\hat{\epsilon}_{\hat{\theta}}(Z^{0}_{t},t,\mathcal{Z}^{\setminus{0}})}{\alpha_{t}}\right),\end{aligned}$
@@
$\begin{aligned} &\mathcal{L}_{\text{CD}}(\hat{\theta},\hat{\theta}^{-};\Psi)=\\ &\mathbb{E}_{n,\mathcal{Z}^{\setminus{0}},Z^{0}_{t_{n+1}}}\left[d\left(f_{\hat{\theta}}\left(Z^{0}_{t_{n+1}},\mathcal{Z}^{\setminus{0}},t_{n+1}\right),f_{\hat{\theta}^{-}}\left(\hat{Z}^{0}_{t_{n}},\mathcal{Z}^{\setminus{0}},t_{n}\right)\right)\right],\end{aligned}$
where $\hat{Z}^{0}_{t_{n}}=Z^{0}_{t_{n+1}}+(t_{n}-t_{n+1})\Psi(Z^{0}_{t_{n+1}},t_{n+1},t_{n},\mathcal{Z}^{\setminus{0}})$ is the solution of the PF-ODE obtained via calling the ODE solver $\Psi$ from time $t_{n+1}$ to $t_{n}$. Here $d(\bm{x},\bm{y})=\|\bm{x}-\bm{y}\|^{2}_{2}$ denotes the discrepancy between two latent points, which penalizes large deviations more heavily and thus provides a smooth, scale-aware objective in the latent space.
IV Experiments